10. There are 10 balls in an urn: 4 red, 3 green, 3 yellow. One randomly takes 3 balls. What is the probability that (a) all of them are of different colors? b) all of them are of the same color? (c) two of them are red? (d) two of them are red and one of them is green?
Answer #1

Ans:

a)P(all different)=4C1*3C1*3C1/10C3

=36/(10*9*8)/(3*2*1)

=216/720=0.3

b)P(all same)=P(all red)+P(all green)+P(all yellow)

=(4C3/10C3)+(3C3/10C3)+(3C3/10C3)

=(4C3+3C3+3C3)/10C3

=(4+1+1)/120=6/120=0.05

c)P(two are red)=4C2*6C1/10C3

=6*6/120=36/120=0.3

d)P(two are red and one is green)=4C2*3C1/10C3

=6*3/120=18/120=0.15